# A High-Speed Hardware Algorithm for Modulus Operation and its Application in Prime Number Calculation

### v1

@@ Abstract
This paper presents a novel high-speed hardware algorithm for the modulus operation for FPGA implementation. The proposed algorithm use only addition, subtraction, logical, and bit shift operations, avoiding the complexities and hardware costs associated with multiplication and division. It demonstrates consistent performance across operand sizes ranging from 32-bit to 2048-bit, addressing scalability challenges in cryptographic applications. Implemented in Verilog HDL and tested on a Xilinx Zynq-7000 family FPGA, the algorithm shows a predictable linear scaling of cycle count with bit length difference (BLD), described by the equation $y=2x+2$, where $y$ represents the cycle count and $x$ represents the BLD. The application of this algorithm in prime number calculation up to 500,000 shows its practical utility and performance advantages. Comprehensive evaluations reveal efficient resource utilization, robust timing performance, and effective power management, making it suitable for high-performance and resource-constrained platforms. The results indicate that the proposed algorithm significantly improves the efficiency of modular arithmetic operations, with potential implications for cryptographic protocols and secure computing.
@@ 1 Introduction
This paper addresses this critical gap by presenting a novel hardware algorithm specifically designed for modulus operation. This focus on the fundamental operation of $A~{}mod~{}B$ distinguishes our work from the majority of studies in the field and offers several key advantages. By optimizing the core modulus operation, our work potentially improves the efficiency of all derived modular operations. This cascading effect could lead to significant performance improvements in systems relying on modular arithmetic.
@@
[…]
@@
    As a practical demonstration, we apply our hardware modulus operation to find prime numbers up to 500,000, showcasing its potential in cryptographic applications.
@@ 2 Related Work
Butler et al. presented a high-speed hardware implementation of the modulus operation optimized for FPGA deployment [15]. They introduced two versions of the algorithm to calculate $x~{}mod~{}z$: one with a fixed modulus and another where the modulus can vary. Their design demonstrated efficient pipelining and resource utilization but faced scalabillity issues for very large operand sizes.
@@ 3 Methodology
3.1 Hardware Algorithm for Modulus Operation
@@
The proposed algorithm computes modulus operation with addition, subtraction, logical, and bit shift operations without using general multiplication or division operations which are expensive in hardware implementation. The pseudocode of the proposed algorithm is given in Figure 1, which calculates $result=A~{}mod~{}B$, where $A$ is the dividend and $B$ is the divisor.
Figure 1: The pseudocode of the proposed hardware modulus algorithm
@@
The algorithm computes the $A~{}mod~{}B$ using state machine approach assuming that $A>B$ and $B>0$. It starts with initializing the state to IDLE and setting variables dividend, divisor, shift, and done to zero. The main loop continuously checks the state. If the state is IDLE and start signal is received, the algorithm assigns $A$ to the dividend and $B$ to the divisor, and the algorithm transitions to the ALIGN state, where the divisor is left-shifted until it aligns with the dividend, incrementing the shift counter at each step.
@@
Once alignment is achieved, the algorithm changes to SUBTRACT state, where it iteratively subtracts the divisor from the dividend if the dividend is greater than or equal to the divisor. The divisor is then right-shifted, and the shift counter is decremented. This process continues until the dividend is smaller than $B$, the shift counter is zero, or the shift counter reaches $N$. Finally, the algorithm changes to the FINISH state, where the result is set to the current dividend, and the state returns to IDLE, indicating the operation is complete. This state machine approach ensures efficient and systematic computation of the modulus operation, making it suitable for FPGA implementation.
@@ 3.1.1 Hardware Implementation of the Modulus Operator
The algorithm was implemented using Verilog Hardware Description Language (HDL). The Xilinx Vivado design tool was employed for simulation and synthesis. Initially, the algorithm was tested with a variety of input values to verify its accuracy in the simulation. Following the initial verification, the Verilog implementation was modified to include the number of cycles consumed for each calculation.
@@
For further analysis, a separate Vivado project was created, instantiating the modulus algorithm to record data directly to the computer. Figure 2 shows the block diagram of the system. The integer values, dividend and divisor, are encoded from the computer and transfered to the First-in-First-Out (FIFO) buffer in the FPGA through 8-bit UART interface. The FPGA system decodes these two integers and send them to the modulus instance. With the start signal modulus block begins its operations and a timer counter also begins to count the clock signals consumed by the modulus block. As soon as the modulus operation completes its operation, done signal is asserted, which indicates to stop the timer counter. Finally, modulus result and the cycle counter values are encoded to be sent to the computer through the FIFO and UART interface. A python program runs on the computer sends and receives the outputs of FPGA system. The recorded data included the dividend, divisor, result, and the number of cycles per calculation. This comprehensive dataset was subsequently analyzed to evaluate the performance and efficiency of the modulus algorithm.
To measure the performance of the algorithm in FPGA hardware, various operand sizes were used. The algorithm can be extended to any bit length of operands by simply changing the integer value $N$ in the algorithm in Figure 1. For this study, operand sizes of 32-bit, 64-bit, 128-bit, 256-bit, 1024-bit, and 2048-bit were considered.
@@ 3.2 Application to Prime Number Calculation
The best way to measure the performance of the hardware algorithm for modulus operation is to use it in a practical application. In this study, we chose prime number calculation, as it involves repeated modulus operations and is computationally intensive. The application calculates all prime numbers up to a given integer. For example, if the given integer is 10, the application finds 2, 3, 5, and 7 as prime numbers. As the given integer increases, the number of computations required grows significantly. Therefore, this application provides an effective means to measure the performance of the new hardware algorithm for modulus operations.
The algorithm in Figure 3 was constructed to identify all prime numbers up to a given integer $A$, using only logical and addition operations. Although this prime number finding algorithm may not be the most optimized solution and there may be better alternatives, it is sufficient for measuring the performance of the new modulus finding algorithm.
@@
Figure 3: Pseudocode for the prime number finding algorithm. This algorithm identifies all prime numbers up to a given integer A serving as a practical application to measure the performance of the new hardware modulus algorithm.
@@ 3.2.1 Hardware Implementation of Prime Number Calculation
Figure 4 shows the datapath diagram of the prime calculation system. The system employs four registers to store the input input integer ($A$), two indexes ($n$ and $i$), and the prime number($P$). The signal $p$ is asserted when $n<A$, and the signal $r$ is asserted when $n$ is equal to 2. Signals $q$ and $t$ are asserted when $i<n$ and $i$ is equal to $n$, respectively. The block mod represents the hardware implementation of the modulus algorithm.
@@
Both hardware implementations, the modulus operation and the prime calculation, were configured using the Xilinx Vivado 2018.3 design suite, targeting a Digilent Zybo Zynq-7000 SoC Trainer FPGA board containing a Xilinx XC7Z010-1CLG400C. Logic resource utilization, power analysis, and timing analysis were conducted using the Vivado design suite for both the hardware modulus implementation and the prime number calculation separately.
@@ 3.3 Measurements
To measure the accuracy and performance of the hardware algorithm for modulus operation, operand sizes of 32-bit, 64-bit, 128-bit, 256-bit, 1024-bit, and 2048-bit were used. Uniform random integers were generated through a Python program and encoded to send to the FPGA board via a UART interface for processing. At the end of each operation, results were sent from the FPGA board to the computer and recorded for further analysis. For each operand size, 10,000 samples were recorded.
@@
To compare the performance of the prime number calculations in FPGA hardware, software implementations of the prime number calculation algorithm (Figure 3) were considered. In the software implementation of the algorithm, the built-in modulus operator (%) was used. Both Python and C programming languages were used to run the same integer set on a Windows 11 PC with 8GB RAM and a processor with a 12th Gen Intel Core i5, 1300 MHz, 10 cores, 12 logical processors, which can operate at a max turbo frequency of 4.9 GHz.
@@ 4 Results and Discussion
4.1 Modulus Algorithm Implementation
Our hardware-implemented modulus algorithm achieved 100% accuracy for all operand sizes when compared against software-based calculations. The Figures 7(a),  7(b), and 7(c) represent the performance of the hardware algorithm for calculating the modulus operation with 32-bit, 64-bit, and 128-bit operands respectively.
@@ 4.1 Modulus Algorithm Implementation
Figure 7: Cycle Count vs Bit Lengths Difference of Operands. Figure 7(a), 7(b), and 7(c) show the experimental data and the linear fit for each operand size. In Figure 7(d) shows the experimental data and the predicted cycle count by $y=2x+2$.
@@
When the BLD is small, regardless of the operand sizes, the calculation completes in very few clock cycles. This is evident from the cluster of data points near the origin of the graph, where both BLD and cycle count are low. This indicates that the proposed hardware algorithm performs very efficiently when the BLD is small.
@@
The Table 2 summarizes the resource utilization of the proposed hardware algorithm implemented on the Zybo FPGA board for different operand sizes (32-bit, 256-bit, 1024-bit, and 2048-bit). The key resources measured include Look-Up Tabes (LUT), Flip-Flops (FF), and Global Clock Buffers (BUFG). The Zybo FPGA has total of LUTs 17,600 and total FFs of 35,200.
@@
Table 2: Resource utilization summary for different operand sizes (32-bit, 256-bit, 1024-bit, and 2048-bit) on the Zybo FPGA.
Implementation
Resource Utilization 32-bit 256-bit 1024-bit 2048-bit
Look Up Tables (LUT) 6 15 24 45
Flip-Flops (FFs) 5 9 11 15
Global Clock Buffers (BUFG) 6 6 6 6
@@
Total Negative Slack (TNS) 0.000 ns
Number of Failing Endpoints 0
Total Number of Endpoints (Setup) 1038
Worst Hold Slack (WHS) 0.049 ns
Total Hold Slack (THS) 0.000 ns
Number of Failing Endpoints (Hold) 0
Total Number of Endpoints (Hold) 1030
Worst Pulse Width Slack (WPWS) 3.500 ns
Total Pulse Width Negative Slack (TPWS) 0.000 ns
Number of Failing Endpoints (Pulse Width) 0
Total Number of Endpoints (Pulse Width) 485
@@
The 2048-bit implementation on the Zybo FPGA board demonstrates robust timing performance, with no violations in setup, hold, or pulse width requirements, and can operate at a maximum frequency of approximately 295.5 MHz. The power consumption is efficiently managed, with a balanced distribution between dynamic and static power components. These results validate the efficacy of the proposed hardware modulus algorithm for high-performance applications on FPGA platforms.
@@
Junction Temperature ${27.0}^{\circ}$ C
Thermal Margin ${58.0}^{\circ}$ C (4.9 W)

### v2

# An Efficient Algorithm for Modulus Operation and Its Hardware Implementation in Prime Number Calculation

@@ Abstract
This paper presents a novel algorithm for the modulus operation for FPGA implementation. The proposed algorithm use only addition, subtraction, logical, and bit shift operations, avoiding the complexities and hardware costs associated with multiplication and division. It demonstrates consistent performance across operand sizes ranging from 32-bit to 2048-bit, addressing scalability challenges in cryptographic applications. Implemented in Verilog HDL and tested on a Xilinx Zynq-7000 family FPGA, the algorithm shows a predictable linear scaling of cycle count with bit length difference (BLD), described by the equation $y=2x+2$, where $y$ represents the cycle count and $x$ represents the BLD. The application of this algorithm in prime number calculation up to 500,000 shows its practical utility and performance advantages. Comprehensive evaluations reveal efficient resource utilization, robust timing performance, and effective power management, making it suitable for high-performance and resource-constrained platforms. The results indicate that the proposed algorithm significantly improves the efficiency of modular arithmetic operations, with potential implications for cryptographic protocols and secure computing.
keywords:
Modular Arithmetic , Hardware Implementation , Modulus Operation , Prime Number Calculation , Cryptographic Applications , FPGA , Verilog HDL
@@
This is the accepted manuscript (post-peer-review, pre-typeset version) of the article published in International Journal of Electronics and Communications (AEÜ)
Citation: W.A. Susantha Wijesinghe, An efficient algorithm for modulus operation and its hardware implementation in prime number calculation, AEU - International Journal of Electronics and Communications, Volume 191, 2025, 155657, ISSN 1434-8411,
DOI: https://doi.org/10.1016/j.aeue.2024.155657
“© 2025. This manuscript version is made available under the CC BY-NC-ND 4.0 license [http://creativecommons.org/licenses/by-nc-nd/4.0/].”
{highlights}
A novel algorithm for modulus operations and its hardware implementation using only addition, subtraction, logical, and shift operations, avoiding complex multiplication and division.
Consistent performance across operand sizes from 32-bit to 2048-bit, with cycle count scaling linearly with bit length difference as y = 2x + 2.
Efficient FPGA implementation achieving 295.5 MHz maximum frequency for 2048-bit operations, with only $7,920$ LUT utilization and 0.174 W power consumption.
Application to prime number calculation demonstrating superior performance over software implementations, particularly for large integers.
High accuracy across operand sizes, with 100% accuracy up to 2048-bit operations, suitable for cryptographic applications.
@@ 1 Introduction
This paper addresses this critical gap by presenting a novel algorithm specifically designed for modulus operation and its hardware implementation. This focus on the fundamental operation of $A~{}mod~{}B$ distinguishes our work from the majority of studies in the field and offers several key advantages. By optimizing the core modulus operation, our work potentially improves the efficiency of all derived modular operations. This cascading effect could lead to significant performance improvements in systems relying on modular arithmetic.
@@
    Our comprehensive evaluation reveals that the cycle count for the operation scales as $y=2x+2$, where $y$ represents the cycle count and $x$ represents the bit length difference of the operands.
@@
    As a practical demonstration, we apply our novel algorithm for modulus operation to find prime numbers up to 500,000, showcasing its potential in cryptographic applications.
@@ 2 Related Work
Butler et al. presented a high-speed hardware implementation of the modulus operation optimized for FPGA deployment [15]. They introduced two versions of the algorithm to calculate $x~{}mod~{}z$: one with a fixed modulus and another where the modulus can vary. Their design demonstrated efficient pipelining and resource utilization but faced scalability issues for very large operand sizes.
@@ 3 Methodology
3.1 Hardware Architecture for Modulus Operation
@@
The proposed algorithm computes modulus operation with addition, subtraction, logical, and bit shift operations without using general multiplication or division operations which are expensive in hardware implementation. The pseudocode of the proposed algorithm is given in Algorithm 1, which calculates $result=A~{}mod~{}B$, where $A$ is the dividend and $B$ is the divisor.
Algorithm 1 The Novel Algorithm for Modulus Operation
1:N-bit integers $A$ and $B$
2:N-bit integer $result$ where $result=\mod(A,B)$
3:Initialize $state\leftarrow IDLE$
4:Initialize $dividend\leftarrow 0$, $divisor\leftarrow 0$
5:Initialize $shift\leftarrow 0$
6:$done\leftarrow 0$
7:while true do
8:    if $state=IDLE$ then
9:         if $start$ then
10:             $dividend\leftarrow A$
11:             $divisor\leftarrow B$
12:             $shift\leftarrow 0$
13:             $state\leftarrow ALIGN$
14:         end if
15:    else if $state=ALIGN$ then
16:         if $divisor\leq dividend$ and $\lnot divisor[N-1]$ and $shift<N$ then
17:             $divisor\leftarrow divisor\ll 1$
18:             $shift\leftarrow shift+1$
19:         else
20:             $state\leftarrow SUBTRACT$
21:         end if
22:    else if $state=SUBTRACT$ then
23:         if $dividend\geq divisor$ then
24:             $dividend\leftarrow dividend-divisor$
25:         end if
26:         $divisor\leftarrow divisor\gg 1$
27:         $shift\leftarrow shift-1$
28:         if $dividend<divisor$ or $shift=0$ or $shift\geq N$ then
29:             $state\leftarrow FINISH$
30:         end if
31:    else if $state=FINISH$ then
32:         $result\leftarrow dividend$
33:         $done\leftarrow 1$
34:         $state\leftarrow IDLE$
35:    end if
36:end while
@@
The algorithm computes the $A~{}mod~{}B$ using state machine approach assuming that $B>0$. This assumption is validated at the system level during input checks before data is fed into the modulus hardware. It starts with initializing the state to IDLE and setting variables dividend, divisor, shift, and done to zero. The main loop continuously checks the state. If the state is IDLE and start signal is received, the algorithm assigns $A$ to the dividend and $B$ to the divisor, and the algorithm transitions to the ALIGN state, where the divisor is left-shifted until it aligns with the dividend, incrementing the shift counter at each step.
@@
Once alignment is achieved, the algorithm changes to SUBTRACT state, where it iteratively subtracts the divisor from the dividend if the dividend is greater than or equal to the divisor. The divisor is then right-shifted, and the shift counter is decremented. This process continues until the dividend is smaller than $B$, the shift counter is zero, or the shift counter reaches $N$. Finally, the algorithm changes to the FINISH state, where the result is set to the current dividend, and the state returns to IDLE, indicating the operation is complete.
3.1.1 Numerical Example
This section provides a detailed example of the the novel algorithm for modulus operation applied to two 6-bit integers, ($A=29$) and ($B=5$), which computes the result of $A\mod B$. This example illustrates how the algorithm operates by tracking changes in key variables: the $dividend$, $divisor$, and $shift$ counter, throughout the execution.
Let $A=29$ and $B=5$ in a 6-bit system $N=6$. The initial values are:
[…]
${Shift}$ $=0$
We will now describe each step of the algorithm.
Align the Divisor:
To begin, the $divisor$ is left-shifted until it is greater than or equal to the $dividend$:
Shift 1: $B\ll 1=10\quad(\text{binary: }00001010_{2}),\quad{Shift}=1$
[…]
At this point, the $divisor$ ( 40 ) is greater than the $dividend$ ( 29 ), so the algorithm transitions to the subtraction phase.
Subtract and Adjust:
In this phase, the $divisor$ is right-shifted, and subtraction is performed when the $divisor$ is smaller than or equal to the $dividend$:
[…]
At the end of this phase, the $dividend$ is 4 and the $divisor$ ( 5 ) is now greater than the $dividend$. Therefore, the algorithm will finish with the remainder.
Final Result:
$result=dividend$
At this point, the $divisor$ has become greater than the $dividend,$ and the $shift$ counter has reached zero. Therefore, the algorithm terminates, and the final result of $29\mod 5$ is $4$.
@@ 3.1.2 Hardware Implementation of the Modulus Operator
The state machine approach ensures efficient and systematic computation of the modulus operation, as shown in Algorithm 1, making it well-suited for FPGA implementation.
Figure 1: Finite State Machine of the modulus algorithm. Here $condition1\equiv(divisor\leq dividend)~{}and~{}(!divisor[N-1])~{}and~{}(shift<N)$ and $condition2\equiv(dividend<B)~{}or~{}(shift==0)~{}or~{}(shift\geq N)$
Figure 1 presents the Finite State Machine (FSM) diagram with four states, operating as follows: Initially, the system remains in the IDLE state until the start signal is asserted. Once the start signal is received, the system transitions to the ALIGN state and stays there until condition1 is met; otherwise, it moves to the SUBT state. In the SUBT state, the system continues until condition2 is met, after which it transitions to the FINISH state, signaling the completion of the calculation. The definitions of condition1 and condition2 are provided in the caption of Figure 1. The operations performed in each state are outlined in Table 1.
Table 1: Operations assigned in each state of the FSM calculating the modulus operation
State Operations
IDLE
$dividend=A$
$divisor=B$
$shift=0$
ALIGN
if $(divisor\leq dividend)~{}and~{}!divisor[N-1]~{}and~{}(shift<N)$:
   $divisor=divisor\ll 1$
   $shift=shift+1$
SUBT
if $dividend\geq divisor$:
   $dividend=dividend-divisor$
$divisor=divisor\gg 1$
$shift=shift-1$
FINISH
$result=dividend$
$done=1$
The Register Transfer Level (RTL) diagram shown in Figure 2 illustrates the operations of the algorithm. Three registers are utilized: one for the dividend, one for the divisor, and one for shifting, which hold the current values. Each register receives its inputs through a separate multiplexer. Additionally, a distinct register is designated to capture the result of the computation. The comparator blocks perform the specified comparisons, while addition, subtraction, and shifting operations are executed on the current values. The outputs of these operations are connected to their respective multiplexers for updating the next values.
Figure 2: Register Transfer Level (RTL) diagram illustrating the hardware architecture for implementing the proposed modulus operation algorithm on an FPGA.
Figure 3: The waveform of the Verilog simulation of the proposed algorithm for modulus operation.
@@
The algorithm for modulus calculation was described using the Verilog Hardware Description Language (HDL) to implement the FSM shown in Figure 1. The Xilinx Vivado design tool was employed for simulation and synthesis. Initially, the algorithm was tested with various input values to verify its accuracy in simulation. Figure 3 shows the input/output waveforms from the simulation of finding (29 mod 5), which is equal to 4.
Following simulations, the Verilog implementation was expanded to include additional modules such as the UART transmitter(uart_tx), UART receiver(uart_rx), cycle counter(cycle_counter), and controller module (Modulus_Controller), as shown in the block diagram in Figure 4.
The operation of the system to calculate $A\mod B$ is as follows: A Python program running on the computer sends the operands—dividend ($A$) and divisor ($B$) as 8-bit data packets through the UART transmitter to the FPGA board. The uart_rx module receives the bitstream through the $rx$ input and outputs 8-bit data to the Modulus_Controller module. The Modulus_Controller module combines the 8-bit data received through uart_rx to form the appropriate bit lengths for operands $A$ and $B$. Afterward, the Modulus_Controller sends $A$ and $B$ to the modulus module and asserts the modulus_start signal. At the same time, the cycle_counter module begins counting clock cycles with the assertion of the modulus_start signal. When the modulus calculation is complete, the modulus module generates the modulus_done signal, and simultaneously, the cycle_counter stops counting. The Modulus_Controller retrieves the result from the modulus module and the cycle count from the cycle_counter, then transmits this information through the uart_tx module to the computer. The Python program running on the computer decodes the relevant data and stores it in a file. The recorded data included the dividend, divisor, result, and the number of cycles per calculation. This comprehensive dataset was subsequently analyzed to evaluate the performance and efficiency of the modulus algorithm.
To measure the performance of the algorithm in FPGA hardware, various operand sizes were used. The algorithm can be extended to any bit length of operands by simply changing the integer value $N$ in the algorithm in Algorithm 1. For this study, operand sizes of 32-bit, 64-bit, 128-bit, 256-bit, 1024-bit, and 2048-bit were considered.
@@ 3.2 Application to Prime Number Calculation
The best way to measure the performance of the novel algorithm for modulus operation is to use it in a practical application. In this study, we chose prime number calculation, as it involves repeated modulus operations and is computationally intensive. The application calculates all prime numbers up to a given integer. For example, if the given integer is 10, the application finds 2, 3, 5, and 7 as prime numbers. As the given integer increases, the number of computations required grows significantly. Therefore, this application provides an effective means to measure the performance of the new algorithm for modulus operations.
The pseudocode in Algorithm 2 was constructed to identify all prime numbers up to a given integer $A$, using only logical and addition operations. Although this prime number finding algorithm may not be the most optimized solution and there may be better alternatives, it is sufficient for measuring the performance of the new modulus finding algorithm.
@@
Algorithm 2 Prime number calculation up to a given integer $A$ using modulus operation
1:Initialize $n\leftarrow 1$, $pCount\leftarrow 0$, $sel_{1}\leftarrow 0$, $sel_{2}\leftarrow 0$, $primes[]$
2:while $n<A$ do
3:    if $n=2$ then
4:         $primes.append(n)$
5:         $pCount\leftarrow pCount+1$
6:    else
7:         $i\leftarrow 3$
8:         while $i\leq n$ do
9:             if $i=n$ then
10:                 $primes.append(n)$
11:                 $pCount\leftarrow pCount+1$
12:                 break
13:             else
14:                 if $sel2=0$ then
15:                     $s\leftarrow n\mod 2$
16:                     $sel2\leftarrow 1$
17:                 else
18:                     $s\leftarrow n\mod i$
19:                     if $s=0$ then
20:                         break
21:                     end if
22:                 end if
23:             end if
24:             $i\leftarrow i+2$
25:         end while
26:    end if
27:    $n\leftarrow n+1$
28:end while
29:return $[pCount,primes]$
@@ 3.2.1 Hardware Implementation of Prime Number Calculation
Figure 5 shows the datapath diagram of the prime calculation system. The system employs four registers to store the input integer ($A$), two indexes ($n$ and $i$), and the prime number($P$). The signal $p$ is asserted when $n<A$, and the signal $r$ is asserted when $n$ is equal to 2. Signals $q$ and $t$ are asserted when $i<n$ and $i$ is equal to $n$, respectively. The block mod represents the hardware implementation of the modulus algorithm.
@@
Both hardware implementations, the modulus operation and the prime calculation, were configured using the Xilinx Vivado 2018.3 design suite, targeting a Digilent Zybo Zynq-7000 SoC Trainer FPGA board containing a Xilinx XC7Z010-1CLG400C. Logic resource utilization, power analysis, and timing analysis were conducted using the Vivado design suite for both the hardware implementation of the novel algorithm for modulus operation and the prime number calculation separately.
@@ 3.3 Measurements
To measure the accuracy and performance of the novel algorithm for modulus operation, operand sizes of 32-bit, 64-bit, 128-bit, 256-bit, 1024-bit, and 2048-bit were used. Uniform random integers were generated through a Python program and encoded to send to the FPGA board via a UART interface for processing. At the end of each operation, results were sent from the FPGA board to the computer and recorded for further analysis. For each operand size, 10,000 samples were recorded.
@@
To compare the performance of the prime number calculations in FPGA hardware, software implementations of the prime number calculation algorithm (Algorithm 2) were considered. In the software implementation of the algorithm, the built-in modulus operator (%) was used. Both Python and C programming languages were used to run the same integer set on a Windows 11 PC with 8GB RAM and a processor with a 12th Gen Intel Core i5, 1300 MHz, 10 cores, 12 logical processors, which can operate at a max turbo frequency of 4.9 GHz.
Figure 8 shows an image of the hardware setup, where the Xilinx Zybo FPGA board is connected to the computer via the JTAG interface. This setup facilitates programming and debugging of the FPGA, while data transfer between the FPGA and the computer is managed through UART communication over the same JTAG interface, enabling real-time testing and performance evaluation.
Figure 8: A picture of the hardware setup, where the Xilinx Zybo FPGA board is connected to the computer via the JTAG interface.
@@ 4 Results and Discussion
4.1 Hardware Implementation of the Novel Algorithm for Modulus Operation
Our hardware-implemented modulus algorithm achieved 100% accuracy for all operand sizes when compared against software-based calculations. The Figures 9(a),  9(b), and 9(c) represent the performance of the hardware implemented algorithm for calculating the modulus operation with 32-bit, 64-bit, and 128-bit operands respectively.
@@ 4.1 Hardware Implementation of the Novel Algorithm for Modulus Operation
Figure 9: Cycle Count vs Bit Lengths Difference of Operands. Figure 9(a), 9(b), and 9(c) show the experimental data and the linear fit for each operand size. In Figure 9(d) shows the experimental data and the predicted cycle count by $y=2x+2$.
@@
When the BLD is small, regardless of the operand sizes, the calculation completes in very few clock cycles. This is evident from the cluster of data points near the origin of the graph, where both BLD and cycle count are low. This indicates that the proposed novel algorithm performs very efficiently when the BLD is small.
@@
The Table 3 summarizes the resource utilization of the proposed algorithm implemented on the Zybo FPGA board for different operand sizes (32-bit, 256-bit, 1024-bit, and 2048-bit). The key resources measured include Look-Up Tabes (LUT) and Flip-Flops (FF).
@@
Table 3: Resource utilization ( Look Up Tables-LUTs and Flip-Flops - FFs) summary for different operand sizes (32-bit, 256-bit, 1024-bit, and 2048-bit) on the Zybo FPGA.
Resources Available Utilization
32-bit 256-bit 1024-bit 2048-bit
LUTs 17,600 1,056(6%) 2,640(15%) 4,224(24%) 7,920(45%)
FFs 35,200 1,760(5%) 3,168(9%) 3,872(11%) 5,280(15%)
@@
Maximum Frequency 295.5 MHz
@@
The 2048-bit implementation on the Zybo FPGA board demonstrates robust timing performance, with no violations in setup, hold, or pulse width requirements, and can operate at a maximum frequency of approximately 295.5 MHz. The power consumption is efficiently managed, with a balanced distribution between dynamic and static power components. These results validate the efficacy of the proposed algorithm for high-performance applications on FPGA platforms.